The diameter of a brand of​ ping-pong balls is approximately normally​ distributed, with a mean of 1.31 inches and a standard deviation of 0.04 inch. A random sample of 4 ​ping-pong balls is selected.

d. The probability is 54​% that the sample mean will be between what two​ values, symmetrically distributed around the population​ mean?

The lower bound is

The upper bound is
Answer #1

z score corresponding to middle 54% area = -0.74, 0.74

Hence,

Lower bound = 1.31 - 0.74*0.04/ = 1.31 - 0.0148 = 1.2952

Upper bound = 1.31 + 0.74*0.04/ = 1.31 + 0.0148 = 1.3248